# Avoid Wasted Annotation Costs in Open-set Active Learning with Pre-trained Vision-Language Model

### v1

@@ 4.2 Open-set Robustness
Figure 11: Changes in OOD detection accuracy of CLIPNAL based on the temperature parameter. Both datasets are constructed with a 40% mismatch ratio and a 60% OOD ratio.
4.3 Ablation Studies
Effect of Visual Similarity Weights.
To improve the accuracy of OOD detection, CLIPNAL uses visual similarity weights from labeled ID data. Fig. 10 shows (a) accuracy and (b) precision improvements with and without visual similarity weights. (a) Accuracy improvement is significant with increasing mismatch ratios and OOD ratios, except for a 20% OOD ratio in CIFAR10. (b) Precision results indicate that visual similarity weights significantly increase the amount of ID data selected, especially in CIFAR100, demonstrating the effectiveness of visual similarity weights in preventing cost loss.
@@ 4.3 Ablation Studies
Necessity of Self-Temperature Tuning.
Fig. 11 shows the accuracy of purity assessment with varying temperature parameters $\tau$. The accuracy of purity assessment changes significantly with small variations in $\tau$, indicating the importance of optimal temperature tuning. The optimal $\tau$ values for 0.108 on CIFAR10 and 0.123 on CIFAR100 differ, emphasizing the need for self-temperature tuning based on labeled data to select high-purity data correctly.
@@ 4.4 Visualization of Query Strategy Results
Fig. 12 visualizes the ID data and informativeness selected by each query strategy. In the case of RANDOM, the selected data shows no specific pattern and is chosen randomly. In contrast, CORESET selects data with diversity, resulting in less redundancy and more varied data than RANDOM. LL and CONF select data based on uncertainty and choose data near the decision boundaries between classes. However, the query strategies in standard AL result in fewer usable training data due to the selection of OOD data. On the other hand, open-set AL methods select more usable training data than standard AL. However, LfOSA and EOAL, focusing on purity rather than decision boundaries, tend to select less informative data. MQNet selects relatively more ID data than LL, but only in certain decision boundary areas. Conversely, CLIPNAL selects data primarily from high informativeness areas near decision boundaries, similar to CONF, and achieves the highest selection of ID data among the query strategies.
@@ 5 Conclusion
This study introduces CLIPNAL, a two-stage query strategy designed to minimize cost losses in open-set active learning by improving the accuracy of purity assessments and selecting highly informative data from ID data only. CLIPNAL utilizes a pre-trained CLIPN model without additional training to detect OOD data based on linguistic information and visual similarity weights derived from collected ID data. To maximize the accuracy of the pre-trained CLIPN’s ID data assessment, CLIPNAL optimizes the temperature parameter through self-temperature tuning based on the labeled ID and OOD data. By accurately selecting ID data from unlabeled data, CLIPNAL can apply a standard AL query strategy to select only highly informative data, reducing cost losses and approximating the conditions assumed in standard AL.
@@
To validate the efficiency of the proposed method, various open-set datasets with different mismatch ratios and OOD ratios were constructed using CIFAR10 and CIFAR100. Experimental results demonstrate that CLIPNAL consistently achieves the lowest cost losses and highest accuracy across all scenarios. As CLIPNAL conducts purity assessment and data informativeness evaluation sequentially, it can incorporate various query strategies proposed in standard AL after assessing purity. This flexibility indicates that future research can extend the application of CLIPNAL to not only classification but also regression, object detection, and object segmentation in open-set scenarios.
Appendix A CLIPNAL Process
Algorithm 1 summarizes the entire CLIPNAL process. AL involves data collection and model training. Initially, since no labeled data exists, data is randomly selected from the unlabeled pool for model training. Given that the unlabeled data contains OOD data, the proportion of OOD data in the selected data reflects the overall distribution. Only ID data is used for model training, excluding OOD data. Once initial training is complete, data is selected using the proposed CLIPNAL method. CLIPNAL assesses the purity of data from the unlabeled pool. First, it explores the optimal temperature parameter using labeled ID and OOD data, then applies CLIPN’s purity assessment process and visual weighting to calculate the purity of the unlabeled data. Based on the calculated purity, OOD data is excluded, and the remaining data is assessed for informativeness using a standard AL query strategy. The selected data is annotated according to predefined costs and added to the labeled data. After data collection, the classifier is trained with the newly labeled data, and the model is reinitialized and retrained upon further data collection. Once the user-defined number of rounds is complete, the final classifier trained with the last labeled data is used as the final model.
1 Input: Labeled data $S=\{y_{i},x_{i}\in S^{\text{ID}}\cup S^{\text{OOD}}\}_{i=1}^{N_{\text{L}}}$, Unlabeled data $D_{\text{U}}=\{x_{i}\in\textbf{x}^{\text{ID}}\cup\textbf{x}^{\text{OOD}}\}_{i=1}^{N_{\text{U}}}$, Query $Q$, classifier $f^{\text{clf}}$, Image encoder $f^{\text{img}}$, Text encoder $f^{\text{text}}$, No text encoder $f^{\text{no text}}$, Temperature $\tau$, Annotation costs $C$, Number of classes in labeled data $K$, and Round $R$.
2Output: $f^{\text{clf}}$
3
4
// Initial labeled data sampling
5 Randomly select samples $S=\{x_{i}\}_{i=1}^{C}$ from $D_{\text{U}}$.
6Divide $S$ into $S^{\text{ID}}$ and $S^{\text{OOD}}$.
7
8for $0\leq r\leq R$ do
9       Initialize $f^{\text{clf}}$
10      Train $f^{\text{clf}}$ on $S^{\text{ID}}$
11      Find optimal temperature $\tau^{*}$ using Eq.8.
12      $\hat{\textbf{x}}_{\text{U}}^{\text{ID}}\leftarrow$ Purity($D_{\text{U}},S^{\text{ID}},f^{\text{img}},f^{\text{text}},f^{\text{no text}},\tau^{*}$)
13      $Q\leftarrow$ StandardAL($f^{\text{clf}},\hat{\textbf{x}}_{\text{U}}^{\text{ID}},C$)
14      Make annotation on the selected $Q$
15      $S^{\text{ID}}\leftarrow S^{\text{ID}}\cup\textbf{x}^{\text{ID}}$
16      $S^{\text{OOD}}\leftarrow S^{\text{OOD}}\cup\textbf{x}^{\text{OOD}}$
17      $D_{\text{U}}\leftarrow D_{\text{U}}-Q$
18return $f^{\text{clf}}$
19
20Function Purity($D_{\text{U}},S_{\text{ID}},f^{\text{img}},f^{\text{text}},f^{\text{no text}},\tau$):
       // Extract the features of unlabeled data
21       Calculate $\textbf{z}^{\text{img}}$ of $D_{\text{U}}$ and $\textbf{z}^{\text{text}}$, $\textbf{z}^{\text{no text}}$ using Eq.1.
      // Calculate ID and OOD probability using ATD
22       Calculate $p_{ij}^{\text{ID}}$ and $p_{ij}^{\text{OOD}}$ with $\tau$, $\textbf{z}^{\text{img}}$, $\textbf{z}^{\text{text}}$, and $\textbf{z}^{\text{no text}}$ using Eq.4
      // Adjust ID and OOD probability using visual similarity weights
23       Calculate class representation $\phi_{j}$ of $S_{\text{ID}}$ using Eq.5.
24      Calculate visual similarity weights $s_{ij}$ with $\tau$ and $\phi_{j}$ using Eq.6.
25      Adjust $p_{ij}^{\text{ID}}$ and $p_{ij}^{\text{OOD}}$ to $p_{ij}^{*,\text{ID}}$ and $p_{ij}^{*,\text{OOD}}$ using 7.
      // Select predicted ID samples
26       Select $\hat{\textbf{x}}_{\text{U}}^{\text{ID}}\leftarrow\{\mathbb{I}^{*}(x_{i})=0,x_{i}\in D_{\text{U}}\}$.
27      return $\hat{\textbf{x}}_{\text{U}}^{\text{ID}}$
28
Algorithm 1 CLIPNAL Process
Figure 13: Example of the data purity assessment process in CLIPNAL.
Appendix B Open-set Data Description
We utilize CIFAR10 and CIFAR100 for our image classification tasks. CIFAR10 and CIFAR100 consist of 10 and 100 classes, each comprising 50,000 training and 10,000 test images. We reconstruct the datasets by employing mismatch and OOD ratios for open-set data. The mismatch ratio represents the proportion of ID classes in the dataset[10, 13], while the OOD ratio denotes the proportion of OOD data in the unlabeled pool[4, 11]. Previous experiments by LfOSA and EOAL focused on open-set data constructed with the mismatch ratio but did not consider the OOD ratio. We incorporate both ratios to create more realistic scenarios when constructing open-set data.
Table 1: The number of samples for open-set scenarios.
Mismatch Ratio Dataset (Class Ratio) OOD Ratio Initial $S_{\text{L}}$ $D_{\text{U}}$
ID OOD ID OOD
20% CIFAR10 (2:8) CIFAR100 (20:80) 10% 450 50 9550 1061
20% 400 100 9600 2400
40% 300 200 9700 6467
60% 200 300 9800 14700
30% CIFAR10 (3:7) CIFAR100 (30:70) 10% 450 50 14550 1617
20% 400 100 14600 3650
40% 300 200 14700 9800
60% 200 300 14800 22200
40% CIFAR10 (4:6) CIFAR100 (40:60) 10% 450 50 19550 2172
20% 400 100 19600 4900
40% 300 200 19700 13133
60% 200 300 19800 29700
Appendix C Implementation Details
The image classifier used in the experiments is ResNet18[6]. The model is trained for 20,000 steps with a batch size of 64 for all datasets. The learning rate was set to 0.1, weight decay to 0.0005, and momentum to 0.9, with the model parameters updated using Stochastic Gradient Descent (SGD). The temperature parameter for purity assessment in CLIPNAL is tuned within the range of 0 to 1 based on the labeled data.
The image encoder and text encoder for CLIPN were pre-trained on LAION-2B[2], and the ”no” text encoder was pre-trained on CC3M[16]. The image encoder uses a ViT-Base model[3] with a patch size of $16\times 16$. For data informativeness, CLIPNAL uses CONF[18], one of the query strategies proposed in standard AL. Hyper-parameters for LL, MQNet, LfOSA, and EOAL are set as per the original studies[21, 11, 10]. Experiments are conducted using Python 3.8.10 and PyTorch 1.14 on an Intel Xeon CPU E5-2698 v4 with a Tesla V100 GPU.
Appendix D Baselines Description
We select query strategies for standard AL based on uncertainty and diversity. Uncertainty-based query strategies include CONF[18], which considers the lowest class prediction confidence of the model for a given data point as uncertainty. If the minimum prediction confidence is high, it implies that the model lacks certainty, thus selecting data with high uncertainty for training. LL[21] adds a Loss Prediction Module to the trained model to predict the loss during training, selecting data points with higher predicted losses as those with higher uncertainty. Diversity-based query strategy: CORESET[15] selects data that encapsulates diverse information to cover the entire dataset, thereby selecting less redundant data containing new information.
We chose query strategies for open-set AL using contrastive learning and OOD detector. Contrastive learning-based query strategy: MQNet[11] uses CSI for OOD detection. It calculates purity using CSI and informativeness using LL, combining the two through meta-learning to define data scores. OOD detector-based query strategy: LfOSA[10] trains a classifier on labeled ID and OOD data to classify $K+1$ classes, then uses a Gaussian Mixture Model to determine data purity. EOAL[13] uses two entropy scores representing purity and informativeness to select data based on these combined scores.
Appendix E Overall Performance on Open-set Data
AL results are assessed using Area Under Budget Curve (AUBC)[22]. AUBC measures the efficiency of AL by calculating the area under the curve of accuracy improvement over rounds. A larger area indicates greater accuracy improvement after data collection.
Table 2 summarizes the AUBC results for all open-set scenarios. MQNet shows high AUBC in CIFAR10 with high OOD and mismatch ratios but does not outperform standard AL in all cases. In CIFAR100, RANDOM often outperforms other strategies. However, CLIPNAL consistently shows higher AUBC than existing methods, indicating significant accuracy improvement from the initial round.
Table 2: AUBC(%) performance of query strategies based on the mismatch and OOD ratio.
Mismatch Ratio Datasets CIFAR10 CIFAR100
OOD Ratio 10% 20% 40% 60% 10% 20% 40% 60%
[…]
CORESET 87.29 87.23 86.49 85.79 62.93 61.57 57.73 52.10
MQNet 86.61 86.56 86.20 85.63 62.68 61.19 56.66 51.19
LfOSA 85.50 85.01 84.57 83.62 58.73 57.46 53.27 47.10
EOAL 86.97 86.59 86.29 85.39 62.70 61.44 56.59 49.62
CLIPNAL 88.10 88.03 87.84 87.64 65.83 65.36 64.01 62.33
[…]
LL 83.78 83.09 81.49 79.50 54.07 51.42 45.38 38.87
CORESET 83.65 83.28 82.13 80.35 53.12 50.88 46.79 40.59
[…]
LfOSA 80.90 80.76 79.59 77.72 48.10 45.93 41.48 36.24
EOAL 83.24 82.94 81.86 79.86 52.40 50.57 45.76 38.13
CLIPNAL 85.50 85.33 85.04 84.40 56.00 55.34 53.75 52.17
[…]
LL 81.47 80.66 78.56 75.50 45.84 43.56 36.94 31.09
[…]
LfOSA 78.13 77.85 76.35 73.56 40.05 38.18 33.86 28.24
EOAL 81.08 80.74 78.93 76.78 44.43 43.22 37.68 30.33
CLIPNAL 83.41 83.20 82.67 81.95 49.52 48.67 47.47 45.65

### v2

@@ 4.3 Ablation Studies
Effect of Visual Similarity Weights. To improve the accuracy of OOD detection, VLPure-AL uses visual similarity weights from labeled ID data. Figure 11 shows (a) accuracy and (b) precision improvements with and without visual similarity weights. (a) Accuracy improvement is significant with increasing mismatch ratios and OOD ratios, except for a 20% OOD ratio in CIFAR10. (b) Precision results indicate that visual similarity weights significantly increase the amount of ID data selected, especially in CIFAR100, demonstrating the effectiveness of visual similarity weights in preventing cost loss.
Table 7: $\text{AUBC}_{acc}$(%) and performance differences (%p) in an ablation study analyzing the effects of Self-Temperature Tuning and Visual Similarity Weights.
CIFAR10 CIFAR100 Tiny-ImageNet Self-Temperature Tuning Visual Similarity weights 10% 60% 10% 60% 10% 60% Best $\text{AUBC}_{acc}$ among baselines 82.38 - 77.15 - 46.20 - 34.27 - 38.02 - 27.28 - Ours 83.41 - 81.95 - 49.52 - 45.65 - 42.67 - 38.72 - 83.37 (-0.04) 81.65 (-0.30) 48.29 (-1.23) 46.11 (+0.46) 41.12 (-1.55) 38.45 (-0.27) 83.17 (-0.24) 81.49 (-0.46) 48.04 (-1.48) 42.46 (-3.19) 40.31 (-2.36) 34.87 (-3.85) 81.22 (-2.19) 79.29 (-2.66) 24.69 (-24.83) 17.29 (-28.36) 17.83 (-24.84) 10.74 (-27.98)
Table 8: Comparison of $\text{AUBC}_{acc}$(%) and performance differences (%p) for different query strategies of standard AL
Datasets CIFAR10 CIFAR100 Tiny-ImageNet OOD Ratio 10% 60% 10% 60% 10% 60% Best $\text{AUBC}_{acc}$ among baselines 82.38 - 77.15 - 46.20 - 34.27 - 38.02 - 27.28 CONF 83.41 (+1.03) 81.95 (+4.80) 49.52 (+3.32) 45.65 (+11.38) 42.67 (+4.65) 38.72 (+11.44) LL 82.25 (-0.13) 80.76 (+3.61) 49.52 (+3.32) 45.77 (+11.50) 40.90 (+2.88) 37.02 (+9.74) VLPure-AL CORESET 81.69 (-0.69) 80.20 (+3.05) 47.48 (+1.28) 44.08 (+9.81) 40.47 (+2.45) 36.81 (+9.53)
Impact of Components. Table 7 presents the ablation study on visual similarity weights and self-temperature tuning. Self-temperature tuning is a method proposed to leverage collected OOD samples effectively. When self-temperature tuning is not applied, the temperature is fixed at 10. The absence of both self-temperature tuning and visual similarity weights leads to a significant performance drop in CIFAR10 and results in poor training outcomes for CIFAR100 and Tiny-ImageNet. This highlights the necessity of applying an optimal temperature. When self-temperature tuning is used to explore the optimal temperature, the method consistently outperforms the baseline across all scenarios. Moreover, even when the optimal temperature is not applied, the inclusion of visual similarity weights achieves better performance than the baseline. This demonstrates that the use of visual similarity weights, derived from labeled data, contributes to improved OOD detection.
Standard AL. VLPure-AL sequentially considers purity first and evaluates informativeness only for samples predicted as ID, enabling the application of various query strategies proposed in standard AL. Table 8 shows the performance differences compared to the baseline when applying three standard AL query strategies: CONF, LL, and CORESET. Except for CIFAR10 with an OOD ratio of 10%, VLPure-AL achieves higher performance than the baseline across all scenarios for all three query strategies.
@@ 4.4 Visualization of Query Strategy Results
Figure 12 visualizes the ID data and informativeness selected by each query strategy. In the case of RANDOM, the selected data shows no specific pattern and is chosen randomly. In contrast, CORESET selects data with diversity, resulting in less redundancy and more varied data than RANDOM. LL and CONF select data based on uncertainty and choose data near the decision boundaries between classes. However, the query strategies in standard AL result in fewer usable training data due to the selection of OOD data. On the other hand, open-set AL methods select more usable training data than standard AL. However, LfOSA and EOAL, focusing on purity rather than decision boundaries, tend to select less informative data. MQNet selects relatively more ID data than LL, but only in certain decision boundary areas. Conversely, VLPure-AL selects data primarily from high informativeness areas near decision boundaries, similar to CONF, and achieves the highest selection of ID data among the query strategies.
@@ 5 Conclusion
This study introduces VLPure-AL, a two-stage query strategy designed to minimize cost losses in open-set active learning by improving the accuracy of purity assessments and selecting highly informative data from ID data only. VLPure-AL utilizes a pre-trained CLIPN model to detect OOD data based on linguistic information and visual similarity weights derived from collected ID data. To maximize the accuracy of the pre-trained CLIPN’s ID data assessment, VLPure-AL optimizes the temperature parameter through self-temperature tuning based on the labeled ID and OOD data. By accurately selecting ID data from unlabeled data, VLPure-AL can apply a standard AL query strategy to select only highly informative data, reducing cost losses and approximating the conditions assumed in standard AL.
@@
To validate the efficiency of the proposed method, various open-set datasets were constructed by applying mismatch ratios and OOD ratios to CIFAR10, CIFAR100, and Tiny-ImageNet. Additionally, diverse experiments were conducted using DomainNet by assuming different domains as OOD. Experimental results demonstrate that VLPure-AL consistently achieves the lowest cost losses and highest accuracy across all scenarios. As VLPure-AL conducts purity assessment and data informativeness evaluation sequentially, it can incorporate various query strategies proposed in standard AL after assessing purity. This flexibility indicates that future research can extend the application of VLPure-AL to not only classification but also regression, object detection, and object segmentation in open-set scenarios.